If Colorado Springs, Colorado, has 1.2 times as many days of sunshine than Boston, Massachusetts, how many days of sunshine does each city have if there are a total of 464 days of sunshine between the two in a year?
Answer by JulietG(1812)

C + B = 464
C = 1.2B
Substitute the known value of C from the second equation into the first
C + B = 464
(1.2B) + B = 464
2.2B = 464
Divide each side by 2.2
B = 210.9, rounded to 211 (it's asking for days, not partial days, so round)
If Boston has 211 days of sunshine (doubtful), then Colorado Springs has 1.2 times that amount. 211 * 1.2 = 253.2 (rounded to 253)
211 + 253 = 464